Suppose a firm's cost function is C (q) = 2q2 + 8.

i) The variable cost is? a) q2 b) 2q2  c) 2 d) 2q

ii) The average fixed cost is? a) 2q2 b) 8 c) 2q d) 8/q

iii) The average total cost is? a)2q2 b) 2q c) 2q + 8/q d) 8/q

iv) The average variable cost is? a) 2q2 b) 2q c) 8 d) 2q + 8/q

vi) The marginal cost is? a) 2q2 b) q2 c) q d) 4q

vii) The fixed cost is ? a) 8 b) 2q2 c) 2 d) q2
Answer #1

In the above cost function Total Fixed Cost(TFC) = 8 and Total Variable Cost(TVC) = 2q2

i) variable cost = 2q2

ii) Average fixed cost = TFC / q = 8 / q

iii) Average total cost = TC / q = (2q2 + 8) / q = 2q + 8/q

iv) Average variable cost = TVC / q = 2q2 / q = 2q

vi) Marginal cost = d(TC) / dq = d(2q2 + 8) / dq = 4q

vii) Fixed cost = 8